OpenStudy (anonymous):

The maximum value of the expression 1/(sin^2x + 3sinx cox + 5 cos^2 x)?

OpenStudy (anonymous):

I guess 1

OpenStudy (anonymous):

the answer should be 2

OpenStudy (jiteshmeghwal9):

\[\frac{1}{(1-\cos^2x)+3sinx \space cosx + 5 \cos^2x}\]

OpenStudy (jiteshmeghwal9):

\[\frac{1}{1+4\cos^2x +3sinx \space cosx}\]

OpenStudy (anonymous):